OpenStudy (vera_ewing):

What is the pH of a solution with [OH-] = 2.6 × 10^-3 M?

OpenStudy (vera_ewing):

@thomaster Help please!

OpenStudy (vera_ewing):

@jcoury help please!

OpenStudy (anonymous):

doing it hold on

OpenStudy (anonymous):

okay i cant figure out how to do it with -3 but this is with -2, you you could just replace the -2 with -3 and do the same steps: HBr is a strong acid so it dissociates completely in water: HBr + H2O = H3O+ + Br- 1mol HBr = 1mol H3O+ [HBr] = 2.6*10^-2. Therefore: [H3O+] = 2.6*10^-2 What is pH? pH = -log [H3O+] pH = -log (2.6*10^-2) pH = 1.59 What is [OH-] pH + pOH = 14.00 pOH = 14.00-pH pOH = 14.00 - 1.59 pOH = 12.41 [OH-] = 10^-pOH [OH-] = 10^-12.41 [OH-] = 3.85*10^-13 There is a very simple way to work out [OH-] [H3O+] * [OH-] = 10^-14 [OH-] = (10^-14) / (2.6*10^-2) [OH-] = 3.85*10^-13 I did both methods because OI am not sure what you have been taught.

OpenStudy (vera_ewing):

Wait but what's the answer?

OpenStudy (anonymous):

3.85*10^-13 But you would have to do the steps i just showed you but with -3

OpenStudy (vera_ewing):

These are my answer choices A. 11.41 B. 5.95 C. 2.59 D. 8.05

OpenStudy (anonymous):

okay hold up. let @yamyam70 take it from here

OpenStudy (vera_ewing):

@yamyam70 Help please!

OpenStudy (yamyam70):

@jcoury @vera_ewing this is chem! Im not good with chem! haha ILL TRY thougn

OpenStudy (vera_ewing):

@Rushwr Please help!

OpenStudy (yamyam70):

@vera_ewing you want to get pH from [ OH- ] ? not pOH right?

OpenStudy (vera_ewing):

Yes, what is the pH?

OpenStudy (yamyam70):

@vera_ewing working with the maths, ill post it here after haha

OpenStudy (yamyam70):

@vera_ewing We need to convert it from [OH-] to [ H+ ] from [H + ] we can get the value of pH by taking the logarithm

OpenStudy (yamyam70):



OpenStudy (yamyam70):

@vera_ewing There you go!

OpenStudy (yamyam70):

@jcoury I found the site where you copy pasted your answer here, that is indeed off topic, and does not help in any form. Please avoid this type of actions. thanks reference https://answers.yahoo.com/question/index?qid=20100809235159AAcCjxE

OpenStudy (anonymous):

I didn't get it from that

OpenStudy (yamyam70):

It is verbatim as it seems.

OpenStudy (vera_ewing):

Thank you so much! :)

OpenStudy (yamyam70):

Welcome :)